### v1

# SLayR: Scene Layout Generation with Rectified Flow

Abstract
We introduce SLayR, Scene Layout Generation with Rectified flow. State-of-the-art text-to-image models achieve impressive results. However, they generate images end-to-end, exposing no fine-grained control over the process. SLayR presents a novel transformer-based rectified flow model for layout generation over a token space that can be decoded into bounding boxes and corresponding labels, which can then be transformed into images using existing models. We show that established metrics for generated images are inconclusive for evaluating their underlying scene layout, and introduce a new benchmark suite, including a carefully designed repeatable human-evaluation procedure that assesses the plausibility and variety of generated layouts. In contrast to previous works, which perform well in either high variety or plausibility, we show that our approach performs well on both of these axes at the same time. It is also at least 5x times smaller in the number of parameters and $37\%$ faster than the baselines. Our complete text-to-image pipeline demonstrates the added benefits of an interpretable and editable intermediate representation.
Figure 1: We introduce SLayR, a method for scene layout generation via rectified flow. SLayR generates editable and interpretable scene layouts, which can be rendered using conditional image diffusion. We demonstrate high plausibility and high variety of our generations.
1 Introduction
Recent advances in text-to-image modeling have focused on training denoising diffusion models to generate images from a prompt encoding and image noise [34, 35, 36, 6, 50]. While these methods provide very impressive results, we demonstrate that they show a lack in the variety of generated scenes when compared with images generated with scene layout conditioning. Furthermore, they lack the interpretability and controllability that a two stage text-to-layout followed by layout-to-image pipeline would natively provide. At the same time, many works have focused on generating images from given layouts [45, 51, 27] but these methods cannot perform general text-to-image generation without a given input layout.
Therefore, we introduce a novel text-to-layout generation model and combine it with an existing conditional image generation model as shown in Figure 1 to form a complete text-to-image pipeline, creating more varied images than the end-to-end image generation baseline with the benefit of a human-interpretable and controllable intermediate representation. Given a CLIP [33] embedding of a global scene prompt, we generate the layout using rectified flow [24], with a modified Diffusion Transformer (DiT) [32]. As baselines, we apply LayoutTransformer [11] with a small adaptation to handle global conditioning, and GPT4o  [29], as large language models (LLMs) also develop abstract representations of scenes and can be used to generate layouts [40, 21, 8]. However, we do not consider LLMs as a direct baseline, as they are trained on different and substantially larger datasets. Finally, we apply InstanceDiffusion [45] to obtain images.
We evaluate the final generated images using traditional metrics, and observe that we achieve state-of-the-art in FID [12] and KID [3] among text-to-layout generation baselines. By performing additional tests on images generated with no or fixed layouts, we find that interestingly, the CMMD [15] and IS [37] metrics indicate high quality on this limited data and thus cannot be considered conclusive for evaluating images that contain multi-object scenes.
To establish a comparison directly on the quality of generated layouts, we introduce novel metrics to measure their plausibility and variety. Additionally, to achieve a reliable assessment, we introduce a repeatable human-evaluation study and use it to validate our findings. We show that our model yields the highest trade-off in generating layouts that are both, diverse and plausible, while being significantly more light-weight and faster than its competitors. Finally, we show that our approach can be conditioned with partial layouts and the native editing capabilities of the two-stage pipeline. In summary, our contributions are as follows:
  • We propose the first model for text-to-layout generation using rectified flow, which can generate complex scene layouts with varying number and types of objects from a generic text prompt, and supports partial layout conditioning.
  • We establish a suite of numerical metrics for generated layout evaluation and a well-designed human-evaluation study that can be used to assess the quality and be repeated by others.
  • We demonstrate that our intermediate layouts result in substantially improved variety of the generated scenes, while having high plausibility at the same time.
  • We demonstrate that our method provides controllability, and that the resulting images can be edited.
2 Related Work
Controlled Text-to-Image Generation. Denoising diffusion modeling initiated a new wave [42, 13, 43] of state-of-the-art text-to-image models  [35, 38]. In practice, it is desirable to have fine-grained user control over generated images that surpasses the capabilities of only conditioning on a single text prompt. One way of achieving this is to structure the latent generation [14, 18, 31], allowing inter- and extrapolation of latents. Our method follows an alternative paradigm of conditional generation [51, 27, 41, 28, 47]. In particular, we consider a conditioning on both a global and local bounding box prompts, following previous works GLIGEN [20] and Instance Diffusion [45]. The input to these methods has to be specified manually though, and therefore our work focuses on actually obtaining these layouts from a simple text prompt. We then leverage the existing InstanceDiffusion [45] approach to transform the generated layout to an image.
Layout Generation. LayoutTransformer  [11] is an autoregressive transformer which completes a partial sequence of object bounding boxes to form an image layout. We argue that a model that generates the layout jointly is a more natural choice, as the task does not provide an intrinsic ordering.
A work that leverages diffusion for layout generation is LayoutDM [4], but this model does not support global prompt conditioning, requiring specification and prediction of bounding boxes for all objects in the scene. LayoutFlow [10] is also related to our work as it generates User Interface (UI) layouts via Flow Matching [23]. However, this model is specialized to the UI layout domain, and thus does not natively support global conditioning. The same applies to other UI layout generators, such as DLT  [19] and LayoutFormer++  [16]. SceneGenie  [7] utilizes SG2Im  [17] to generate a scene layout from a pre-existing scene graph. As we assume only a simple global prompt, it is not applicable to our setting.
LLMs in Layout Generation. LLMs have been shown to address related layout generation tasks. LLM-grounded Diffusion [21] takes a multi-object prompt and generates a scene layout, with a bounding box for each object in the prompt. LayoutGPT [8] addresses this task as well, but using a different prompt template. Other works have demonstrated that LLMs can also reason about creating 3D layouts [53, 9, 49, 54, 1]. However, we do not focus on 3D scene generation and consider these works out of scope.
To reasonably measure an LLM’s aptitude for our task, we prompt GPT4o using our own template written to encourage chain of thought reasoning [46]. We measure against GPT4o’s performance in the layout generation task, although it was trained on significantly more data and cannot be considered a direct baseline.
Figure 2: Method Overview. Our layout generation model takes a set of noisy tokens and a global PCA-reduced CLIP embedding as input and uses a transformer-based rectified flow architecture to denoise the tokens into a scene layout that can then be transformed into images with InstanceDiffusion [45]. In order to decode the PCA-reduced CLIP embeddings back to be interpreted by humans, we compare them with a bank of predefined labels. The explicit intermediate layout representation of our pipeline allows editing the scene.
Rectified Flow. Diffusion modeling has inspired numerous variants and improvements, one of which is rectified flow [24]. Prior works on the related text-to-image generation task [25, 6] have successfully leveraged it in their pipelines, with the added benefit of not requiring a noise schedule like DDIM [43] or DDPM [13]-based approaches. This motivates us to use rectified flow for our model.
3 Method
The central part of our work is the text-to-layout generation module, which we combine with the existing InstanceDiffusion [45] to form a complete text-to-image pipeline. An overview is provided in Fig. 2.
3.1 Rectified Flow Preliminaries
We briefly recap the basics of rectified flow introduced in  [24]. Let $I$ be a set of training sample indices and $\{\mathbf{x}_{i}\}_{i\in I}$ the ground-truth samples whose distribution we would like to learn using our model $v$. We linearly interpolate between Gaussian noise $\mathbf{x}_{i}(0)$ and samples $\mathbf{x}_{i}(1)\equiv\mathbf{x}_{i}$ across timesteps $t\in[0,1]$ as follows:
$\mathbf{x}_{i}(t)=(1-t)\cdot\mathbf{x}_{i}(0)+t\cdot\mathbf{x}_{i}(1)\,.$ (1)
The model $v$ is trained to take $(\mathbf{x}_{i}(t),t)$ as input and predict the derivative of the path between $\mathbf{x}_{i}(0)$ and $\mathbf{x}_{i}(1)$, which according to Equation 1 is $\mathbf{x}_{i}(1)-\mathbf{x}_{i}(0)$. The model is trained with the following objective:
[…]
using stochastic gradient descent. This optimization is carried out across all available samples of the ground-truth distribution. Following [24], noisy values $\mathbf{x}_{i}(0)$ are resampled at each epoch. The end result is a network $v$ which is effective at predicting the direction from a noisy sample at an intermediate timestep towards the target distribution. Since a single evaluation may be noisy, the inference is performed by integrating over $T$ timesteps:
$\mathbf{x}_{i}(1)=\mathbf{x}_{i}(0)+\sum_{t=1}^{T}v(x((t-1)/{T}),t/T)\cdot\frac{1}{T}\,.$ (3)
3.2 Layout Representation
We start with defining a scene representation as the basis for our generative architecture. A training sample $(\mathbf{x}_{i},P_{i})$ is composed of a global image caption prompt $P_{i}$ and a set of $J$ object tokens $\mathbf{x}_{i}=\{\mathbf{x}_{i}^{j}\in\mathbb{R}^{d+5}\}_{j\in J}$. The token representation of any single object in the scene is composed of
$\mathbf{x}_{i}^{j}=({\mathbf{b}_{i}^{j}}\mathbin{\|}{\mathbf{c}_{i}^{j}}\mathbin{\|}\alpha_{i}^{j}),$ (4)
where ${\mathbf{b}_{i}^{j}}=(x,y,w,h)\in\mathbb{R}^{4}$ encodes the bounding box coordinates, ${\mathbf{c}_{i}^{j}}\in\mathbb{R}^{d}$ is a PCA-reduced CLIP [33] embedding, and $\alpha_{i}^{j}\in\mathbb{R}$ is an opacity value that defines the existence of a specific bounding box.
3.3 Our Model Architecture
Our rectified flow model is built from multihead AdaLN transformer blocks, which can process tokens $\{\mathbf{x}_{i}^{j}\}_{j\in J}$ to iteratively denoise them [32]. The timestep $t$, bounding box coordinates $\mathbf{b}_{i}^{j}(t)$, and opacity values $\alpha_{i}^{j}(t)$ are sinusoidally encoded. The timestep $t$ and a linear projection of the global label $P_{i}$’s CLIP encoding are passed as conditioning of the adaptive layer normalization of the transformer blocks. The tokens represent the objects in the layout and are processed all at once.
Inference begins at $t=0$ with the set of tokens $\{\mathbf{x}_{i}^{j}(t)\}_{j\in J}\equiv\{\mathbf{x}_{i}^{j}(1)\}_{j\in J}$ initialized from Gaussian noise. Our model then iteratively processes and updates the tokens from $t=0$ to $t=1$ over $T$ iterations using Eq. 3 based on the global prompt conditioning $P_{i}$. As the output of the transformer blocks represents the rate of change of $\mathbf{x}_{i}^{j}(t)$, we use a linear layer to downproject the output back to the dimension of $\mathbf{x}_{i}^{j}(t)$ before sinusoidal encoding.
Using Eq. 3 to sum over all timesteps yields the final layout $\{\mathbf{x}_{i}^{j}(1)\}_{j\in J}$ with PCA-reduced CLIP embeddings, bounding boxes, and opacities. Tokens with $\alpha_{i}^{j}(1)<0.5$ are considered unused and discarded. To determine object labels from the produced tokens, we unproject each $\mathbf{c}_{i}^{j}(1)$ from the PCA space into the CLIP feature space, and find the nearest CLIP embedding from a bank to produce interpretable labels. For image generation, we pass the unprojected embeddings directly into the downstream image generation module.
We explain how we adapt the RePaint [26, 39] technique for rectified flow to enable partial layout conditioning in the supplemental material. As we show in Sec. 5.8, our model can accommodate partial layouts where only some boxes or some labels are given.
3.4 Training
To construct a training sample from the ground-truth image layout $i$, we create ${\mathbf{c}_{i}^{j}}$ and ${\mathbf{b}_{i}^{j}}$ for each bounding box $j$, and initialize $\alpha_{i}^{j}:=1$. To ensure a consistent amount of tokens, we pad the samples by adding tokens with $\alpha_{i}^{j}:=0$, and ${\mathbf{c}_{i}^{j}}$ and ${\mathbf{b}_{i}^{j}}$ sampled from Gaussian noise.
We now treat $\{\mathbf{x}_{i}^{j}\}_{j\in J}\equiv\{\mathbf{x}_{i}^{j}(1)\}_{j\in J}$, sample $\{\mathbf{x}_{i}^{j}(0)\}_{j\in J}$ from Gaussian noise, draw $t$ uniformly from $[0,1]$, and compute the set of tokens $\{\mathbf{x}_{i}^{j}(t)\}_{j\in J}$ by adapting the formula from  Eq. 1, which are then passed to the model as input. We refer to the output of the model as $v(\{\mathbf{x}_{i}^{j}(t)\}_{j\in J},t,P_{i})$ and compute the training loss derived from Eq. 2:
$\mathcal{L}=\sum_{i\in I,j\in J}||\mathbf{x}_{i}^{j}(1)-\mathbf{x}_{i}^{j}(0)-v(\{\mathbf{x}_{i}^{j}(t)\}_{j\in J},t,P_{i})_{j}||^{2}\,.$ (5)
4 Metrics
Assessing the quality of our generated layouts is challenging. Solely evaluating the final images is not meaningful, as their quality is heavily impacted by the choice of conditional image generator and in addition we reveal in Sec. 5.5 that existing metrics are not conclusive. In the following, we therefore contribute novel suitable metrics that allow a reliable evaluation.
4.1 Generated Layout Numerical Metrics
To assess the quality of the generated layout distribution, we introduce four metrics that compare the plausibility and the variety of generated objects.
Object Numeracy.
Our metric $O_{\text{Num}}$ assesses whether generated layouts contain the objects at the expected frequencies. We sample across a collection of global prompts ($\{P_{i}\}$). The probability distribution for expected occurrences of the object-label $\ell$ in layouts generated from the global prompt $P_{i}$ is written $q_{i}^{\ell}$, and the ground-truth probability distribution is $p_{i}^{\ell}$. Our metric is the normalized sum of KL-divergence between these two distributions:
$O_{\text{Num}}:=\frac{\sum_{i,\ell}KL(p_{i}^{\ell}||q_{i}^{\ell})}{|\{P_{i}\}|}\,$ (6)
where lower scores indicate that the model produces layouts with more plausible object numeracy.
Positional Likelihoods.
We introduce $l_{\text{Pos}}^{(1)}$, and $l_{\text{Pos}}^{(2)}$ to measure how plausible the objects in a generated layout are arranged. Let $m$ index all bounding boxes of object-label $\ell$ for prompt $i$. For each object-label $\ell$, we obtain a distribution $k_{i}^{\ell}$ with KDE of the object’s bounding box $({\mathbf{b}_{i}^{\ell}})_{m}$ in all layouts with global prompt $i$. We compute the average likelihood over all objects and all global prompts, to measure the first order positional likelihood:
$l_{\text{Pos}}^{(1)}=\frac{\sum_{i,\ell,m}k_{i}^{\ell}(({\mathbf{b}_{i}^{\ell}})_{m})}{|\{({\mathbf{b}_{i}^{\ell}})_{m}\}|}\,.$ (7)
A higher value for $l_{\text{Pos}}^{(1)}$ means that object bounding boxes are placed in reasonable locations in the layout.
We also want to measure the likelihood of spatial relationships between objects. Let $m^{*}$ index all bounding boxes of object-label ${\ell}^{\prime}$ For each object-labels pair $(\ell,{\ell}^{\prime})$, we obtain a distribution estimated with KDE $k_{i}^{{\ell},{{\ell}^{\prime}}}$ for the difference in the bounding box dimensions. We compute the average likelihood over all objects and all global prompts from our distributions to measure the second order positional likelihood:
$l_{\text{Pos}}^{(2)}=\frac{\sum_{i,\ell\neq{\ell}^{\prime},m,m^{*}}k_{i}^{{\ell},{{\ell}^{\prime}}}(({\mathbf{b}_{i}^{\ell}})_{m}-({\mathbf{b}_{i}^{{\ell}^{\prime}}})_{m^{*}})}{|\{({\mathbf{b}_{i}^{\ell}})_{m}\}|(|\{({\mathbf{b}_{i}^{\ell}})_{m}\}|+1)/2}\,.$ (8)
A higher value for $l_{\text{Pos}}^{(2)}$ means that pairs of objects are plausibly positioned relative to one another. We conduct a grid search across bandwidths with 5-fold cross validation to optimize the KDE bandwidths for both $l_{\text{Pos}}^{(1)}$ and $l_{\text{Pos}}^{(2)}$.
Positional Variance.
Our metric $\sigma^{2}_{\text{Pos}}$ measures the variety of bounding boxes. For every bounding box $({\mathbf{b}_{i}^{\ell}})_{m}$, we find the bounding box in layouts with global prompt $i$ and object label $\ell$ that is closest in Euclidean distance to the bounding box. We now redefine $\{m^{*}\}$ as the set of indices of bounding boxes in other samples which minimize the term $||({\mathbf{b}_{i}^{\ell}})_{m}-({\mathbf{b}_{i}^{{\ell}^{\prime}}})_{m^{*}}||$. We compute all of these Euclidean distances and take the average:
$\sigma^{2}_{\text{Pos}}=\frac{\sum_{i,\ell,m}\sum_{\{m^{*}\}}||({\mathbf{b}_{i}^{\ell}})_{m}-({\mathbf{b}_{i}^{{\ell}^{\prime}}})_{m^{*}}||}{\sum_{i,\ell,m}|\{m^{*}\}|}\,$ (9)
If this metric is small, this means that the variance is low.
4.2 Human Evaluation
Previous metrics capture whether objects appear in similar frequencies, positions, and stand in pairwise relationships similar to the ground-truth distribution. Truly assessing a generated layout requires full spatial and semantic interpretation that can only be carried out by humans. To this end, we introduce a human-evaluation study to assess the plausibility and variety of the final generated images.
To conduct the evaluation, we survey $60$ participants by presenting them images from different methods. For each method, each participant rates $10$ collections for their plausibility, and different $10$ collections for their variety, on a Likert scale ($1$ to $5$, where $1$ corresponds to very implausible/very low variance, and $5$ corresponds to very plausible/very high variance). For plausibility, we instructed participants to consider the overall realism of the collection, as well as how effectively it depicts the global text prompt. For variety, we instructed users to consider the spatial arrangement of objects in an image and implied camera angle in addition to overall image appearance.
A single evaluation of the study on Connect CloudResearch costs about \$160 and can be repeated by others. We provide the full details of the study in the supplemental material. The study was approved by the Ethics Review Board of our institution, ensuring compliance with ethical standards.
5 Experiments
5.1 Training and Evaluation Data
We use the ADE20K dataset [52] that contains approximately 27K images and manual semantic segmentations for indoor and outdoor scenes and a rich collection of object arrangements. The captions reflect the scene category with no additional detail (e.g., “beach”, “lecture room”). From captions and the semantic segmentations, we derive our ground-truth layouts. We use only the top $30$ largest bounding boxes per sample, as this is the default maximum number of bounding boxes supported by InstanceDiffusion  [45], padding samples with fewer bounding boxes.
5.2 Baseline Layout Generation Models
LayoutTransformer. We adapt LayoutTransformer  [11] to our task by training it on our dataset, and enabling it to take a global conditioning as input. To this end, we supply a bounding box that covers the entire image, with the attached global prompt as the first token.
GPT4o. LLMs have been used in related layout generation tasks successfully. Since they have a much larger size and are trained on much more different data, we do not consider them a direct baseline, but interesting to compare to. We selected GPT4o due to its effective performance across a wide range of tasks  [29], and since it was also trained on images. Querying the model with prompts from prior work  [21, 8] yielded empty layouts in $2/3$ of cases, as these templates assume details of the objects to be placed, e.g. “a park with a bench and a tree”. Therefore, we adapted the LLM-grounded Diffusion [21] template that provides seven in-context examples and instructs the model to decompose a generic scene prompt into subcomponents by utilizing chain of thought (COT) reasoning  [46]. We adopt the layout notation format and hyperparameters from  [21].
No Layout Structure. To understand the effect of layout conditioning on image generation, we also run our image generation pipeline with a layout containing no bounding boxes and just a global prompt.
Fixed Layout. To create an adversarial dataset with nonsensical layouts, we also create data with a fixed layout by distributing three bounding boxes evenly-spaced horizontally across the image and randomly sampling labels for each box.
5.3 Layout and Image Sampling
ADE20k is a relatively small dataset, and the training data from classes with few samples can be too little. To reduce the effect from these classes, we use the $15$ highest represented categories and add in five randomly selected categories. For each evaluated model, we generate $30$ layouts for all $20$ selected prompts, and an image conditioned on each layout and corresponding global prompt. This collection of images is also feasible to assess with our human evaluation.
5.4 Visual Results
We provide a qualitative overview of the generated layouts and the final images  Fig. 3. We label bounding boxes using the technique described in  Sec. 3.3, with all object labels present in ADE20k as our word bank.
LayoutTransformer struggles with arranging objects spatially, leading to images with unclear structure that do not depict the global prompt. GPT4o creates plausible layouts, but it struggles in making a variety of arrangements and representing the distribution of possible layouts. Our method produces the most plausible and diverse images across a range of global prompts, including both indoor and outdoor scenes.
“Living Room” “Roof Garden” “Street”
No Layout Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
GPT4o Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
LayoutTransformer Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Ours Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 3: Qualitative comparison (Best viewed up close). Layout objects that are depicted in the generated image are highlighted and labeled. No Layout produces scenes of little variation in viewpoint. GPT4o’s layouts are sensible, but also lack in variety. Layout Transformer produces layouts with implausible arrangements of objects, leading to images which do not depict the global prompt accurately. Our method creates the most plausible and variable layouts, leading to images that are diverse and look realistic.
Model FID ($\downarrow$) KID ($10^{-2}$)($\downarrow$) CMMD ($\downarrow$) IS ($\uparrow$) Inference Time (s) ($\downarrow$) Added Parameters ($\downarrow$)
LayoutTransformer 0.440.440.440.440.44 0.940.940.940.940.94 $1.34$ 8.348.348.348.348.34 24.6124.6124.6124.6124.61 23M
GPT4o $0.94$ $0.99$ $1.34$ $7.87$ $110.69$ $>$1T
Fixed Layout $1.16$ $2.31$ 1.321.321.321.321.32 $8.08$ N/A N/A
No Layout $5.06$ $1.11$DeclareFontSeriesDefault[rm]bfb1.001.001.$0$DeclareFontSeriesDefault[rm]bfb8.508.508.508.50 N/A N/A
[…]
Table 1: Traditional Metric Comparison. Our method performs best on FID and KID while images generated from no layout, while images generated with a fixed layout perform second-best on CMMD, and third-best on IS, but both performed poorly on FID and KID. This suggests that FID and KID are more reliable metrics for the quality of the generated conditioning layout, and that CMMD and IS are not conclusive. Out of all the effective layout generation methods, ours is also the fastest and introduces the fewest parameters.
Partial Conditioning Generated Layout Generated Image Partial Conditioning Generated Layout Generated Image
Refer to caption Refer to caption Refer to caption The scene should contain these objects:   ceiling   wall (2x)   curtain   sink   floor   bathtub   windowpane Refer to caption Refer to caption
Figure 4: Disentangled Generation. We show examples of two modes of disentangled generation for scenes with the prompt Snowy Mountain, and Bathroom, respectively. As visible, our model is able to complete the initial layouts well.
5.5 Generated Image, Size, and Speed Metrics
To evaluate the generated images, we apply the established metrics FID [12], KID  [3], CMMD [15], and IS [37]. While FID, KID, and CMMD compare the distribution of generated images with the ground-truth ones, IS measures the variety of the generated images combined with how well the images can be classified as an object from a fixed set. Since the conditioned image generator may itself introduce details that alter the layout, and we want to limit these effects, we establish the ground-truth images by running InstanceDiffusion on the ground-truth layouts.
To compare the runtimes, we measure the average time it takes for each model to generate a batch of $30$ layout samples on an Nvidia A6000 GPU. Our model and LayoutTransformer generate these samples in parallel. We do not have direct access to GPT4o, and can only query it sequentially through an API. We also record the approximate number of model parameters that our baselines add to a complete text-to-image pipeline, factoring out the weights required for InstanceDiffusion.
We provide a quantitative overview in  Tab. 1. Our method performs state-of-the-art on FID and KID scores. The No Layout and Fixed Layout perform the poorest on these metrics, which indicates that these metrics are robust in effectively measuring structural similarity to the ground-truth images. We observe that the IS and CMMD metrics were optimal when there was no underlying structure at all, with our adversarial Fixed Layout performing second-best on CMMD, and third-best on IS. This indicates that CMMD and IS are ineffective in measuring the effect of the conditioning structure. Because CMMD is a metric which measures the similarity of images in the CLIP feature space, and is meant to align with a text encoding of a corresponding image caption, this metric arguably summarizes the images into global prompts and tends to be agnostic to the image’s spatial structure. In addition, using No Layout will effectively create an image with semantic matching of the global prompt and undermines CMMD’s effectiveness in assessing the underlying layout.
While IS is often used in assessing image generation techniques, it has been criticized for being ineffective at measuring image quality  [2, 5]. Images with multi-object structure might be at a disadvantage, as it measures how well the image fits into one of the single-object classification categories of the Inception-v3 model [44]. Furthermore, while IS is used for judging the variety of generated images, we show in  Sec. 5.7 that our method creates more varied images than the methods which are performing well on IS. These inconclusive findings led us to construct a user study. It is important to note that out of the effective layout generation methods we tested ours has the lowest number of parameters by over a factor of $5$, and is $37\%$ faster than the next fastest competitor.
5.6 Numerical Metrics Results
We show the numerical results from  Sec. 4.1 in  Fig. 5. LayoutTransformer performs comparable to our method in the object numeracy score ($O_{\text{Num}}$) and outperforms us in the position variance, meaning that it produces a large variety ($\sigma^{2}_{\text{Pos}}$) of the right types of objects. However, Figure 3 clearly shows that the generated layouts are not plausible.
GPT4o performs poorly on $O_{\text{Num}}$, although it was likely trained on much more data. When we examined the layouts, we observe that it tends to remain within a much stricter set of possible object labels than our method or LayoutTransformer. While it does well on $l_{\text{Pos}}^{(1)}$, it does poorer than our method on $l_{\text{Pos}}^{(2)}$, indicating that it has a reasonable understanding of where objects should be placed in the scene, but falls short in modeling the correct object interactions. Finally, our Fixed Layouts perform poorly as expected, which demonstrates the accuracy of the metrics.
[…]
[width=]sections/results/figures/baseline_bar_graph_combined
Figure 5: Quantitative Comparison. A comparison of our metrics described in Sec. 4.1 on our method and the layout-generating baselines introduced in Sec. 5.2. Our method is close or outperforms the baselines on $O_{\text{Num}}$, $l_{\text{Pos}}^{(1)}$, and $l_{\text{Pos}}^{(2)}$. Although we are lower than LayoutTransformer on $\sigma^{2}_{\text{Pos}}$,  Fig. 3 clearly shows that the generated variety from Layout Transformer is not plausible. This is further supported by our human evaluation in  Sec. 5.7.
5.7 Human Evaluation
We show the results from our human evaluation study in  Fig. 6. Aligned with the observations from Fig. 3, LayoutTransformer produces the highest variety, but the generated images are the least plausible. GPT4o delivers a similar plausibility as our method but is significantly more limited in variety, while No Layout is the most plausible but least diverse. The latter reveals the important insight that pure text-to-image generative denoising diffusion models do capture significant variety, and that pipelines with intermediate explicit representations can outperform them. We observe that our method really provides the highest trade-off in plausibility and variety, indicating that it can accurately generate a diverse distribution of images, in addition to being much smaller than its competitors.
Figure 6: Human Survey Results. Our method offers a superior trade-off between plausibility and variety, while even being much smaller. The error bars indicate the standard errors of the means and show that the results are statistically significant.
5.8 Additional Model Features
Finally, we present additional features which our model offers, which make it appealing for use in an end-to-end image generation pipeline. In  Fig. 4, we show examples of our model’s performance in different partial layout generation settings. This feature gives users even more fine-grained control over the image generation process. Furthermore, we demonstrate how our pipeline allows for editing of generated images in  Fig. 7 without introducing any additional model components. This is accomplished through modifying the intermediate scene layout, and rerunning InstanceDiffusion with the original seed and global prompt. We observe that the edits are not totally isolated to the altered objects and leave it for future work to develop improved conditional diffusion models.
User Action Image Layout Generated Image
[…]
The “plant” is moved. The generated image is faithful to the new layout. Refer to caption Refer to caption
The “plant” bounding box is removed entirely. The plant is also removed in the generated image. Refer to caption Refer to caption
[…]
Figure 7: Editing. We show how our pipeline enables user editing of images by altering the intermediate scene layout representation. Individual objects can be easily moved, removed, and replaced.
6 Conclusion
We have introduced a text-to-image generation pipeline with an intermediate human-interpretable and controllable layout representation. While other methods produce images of scenes that are either very implausible or of little variety, our method clearly performs well on both of these axes and improves upon the state of the art, while being substantially smaller and faster. Our results show that that plain end-to-end pipelines are limited in the variety that they can capture and that an interpretable intermediate representation can capture this variety much better, while maintaining a high plausibility. In addition, we have introduced a suite of metrics for image layouts and generated images that allow reliable evaluation and can be repeated by the community. Both of these contributions together give rise to exploring image generation pipelines with intermediate explicit representations more in the future. In further work, we will scale up the approach to larger training datasets and add more detailed intermediate phases with segmentation masks.
7 Acknowledgements
We thank Tom Fischer, Raza Yunus, and Julian Chibane for their feedback on the draft. Funded in part by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – GRK 2853/1 “Neuroexplicit Models of Language, Vision, and Action” - project number 471607914.
Supplementary Material
The supplementary material is structured as follows. First, we present the full details of the human evaluation study performed to judge the generation quality in Sec.  A. Next, we introduce details about our partial conditioning procedure in Sec. B. We further provide detailed information about training data and hyperparameters in Sec. C, an ablation study with different of image generators in Sec. D, discussions about LayoutTransformer and GPT4o temperatures in Sec. E and F, respectively, and the GPT4o query template in Sec. G. Last, we provide a comparison between flow matching and DDIM in Sec. H and high-resolution results in Sec. I and Sec. J.
Appendix A Human Evaluation Details
Study Goal. Given the inconclusive results of the traditional metrics from Tab. 1, we aim to establish an effective metric for the quality of the generated layouts with respect to their effect on the final images. For the most accurate assessment, we provide a human-evaluation study that can be repeated by others.
In general, we expect that our generator should generate layouts that appear plausible and are also of a large variety. However, assessing human opinions for these criteria directly on layouts is challenging: the evaluators require time to understand the layout diagrams and explain them, and furthermore, assessments are hard to make without actually seeing the image. To make the study effective, we measure the effect of our model on the downstream generated images. Image qualities that are assessed in other studies (for example the overall quality and aesthetic appeal of the image in Liang et al. [22]) are highly dependent on the conditioned image generator. Therefore we consider these misleading for our case and introduce a suitable study in the following.
Design Principles. We follow the design principles presented by Otani et al. [30] in their work on human evaluation of text-to-image generation: 1) the (evaluation) task should be simple, and 2) results should be interpretable. Following these principles, we show participants only images, and omit the underlying image layouts entirely, which may take some effort to understand. To make the results interpretable, participants rate these images for their plausibility and variety on a Likert scale (as specifically recommended in Otani et al. [30]) from $1$ to $5$. This way, average ratings for different layout generation models can be meaningfully compared to each other, which would be more difficult in other systems (e.g. using non-numbered ranking). To ensure cost efficiency, our survey must be small enough that the data can be collected quickly and repeatedly throughout the model development, and thus we show participants collections of images rather than individual pairs. We kept all of these constraints in mind when designing our study, which is explained in further detail below.
Study Description. Our study was developed using Qualtrics, a standard survey platform. Each participant answers ten plausibility questions and ten variety questions, meaning they rate $80$ image collections in total. We survey $60$ participants. The prompts, image collection index, and the order in which the collections are displayed to participants is randomized to control for any potential effects of a fixed ordering.
We assess the plausibility and variety of generated images that are conditioned on layouts from our method and three of the baselines presented in  Sec. 5.2: LayoutTransformer, GPT4o, and No Layout. We do not test against Fixed Layout, because this was constructed solely as an adversarial example for traditional metrics.
Survey data is selected as described in Sec. 5.3. As shown in Fig. 9, each survey question shows collections of three images from each of the four methods listed above, where every image on the screen has the same global prompt. Given the instructions from Fig. 8, the participant must rate each collection for either their plausibility or variety. Ratings are on a Likert scale ($1$ to $5$, where $1$ corresponds to very implausible/very low variance, and $5$ corresponds to very plausible/very high variance). For plausibility, we instructed participants to consider the overall realism of the collection, as well as how effectively it depicts the global text prompt. For variety, we instructed users to consider the spatial arrangement of objects in an image and implied camera angle in addition to overall image appearance.
Participant Selection and Ethics. Participants were recruited through Connect CloudResearch, a crowdsourcing service built on Amazon Mechanical Turk that implements rigorous quality control procedures to enhance the reliability of the participant pool in line with the study recommendations given by Otani et al. [30]. The study was approved by the Ethics Review Board of our institution, ensuring compliance with ethical standards. Prior to engaging in the tasks, all participants provided informed consent. The study was designed to be completed within $10$ minutes by each participant, who were compensated at an hourly rate of $13.02$ USD. This results in a total cost of 163 USD per run. Participants were anonymized, and we did not collect any personally-identifiable information.
Section 1: Plausibility
“For the following section of the survey, you will be asked to rate collections of images based on how plausible they appear to be, from very implausible to very plausible. An image is considered plausible if objects within the image are realistically and organically placed, and it is a reasonable match to the presented caption. The images do not have to be photorealistic to be considered plausible. You will perform ratings on 10 categories of images, and each page will contain 4 collections that you must rate separately.”
Section 2: Variety
“For the following section of the survey, you will be asked to rate collections of images based on their perceived variance, from very low variance to very high variance. When judging the variance, consider criteria such as the differences in the spatial arrangement of objects, the differences in camera perspective, and the differences in the overall image appearance across the collection. You will perform ratings on 10 categories of images, and each page will contain 4 collections that you must rate separately.”
Figure 8: Full instructions to participants for both sections of the survey. Our instructions clearly define the task and give users detailed information on what to assess
Figure 9: An example question page from our survey. Users must rate collections of 3 images from very implausible to very plausible. The underlying layout generators for the collections shown are (from top to bottom): No Layout, GPT4o, LayoutTransformer, and our method. Collection order was randomized for each question presented to the participant. Users click the button
Appendix B Implementation of Partial Conditioning
We explain our adaptation of the RePaint [26] technique mentioned in  Sec. 3.3, which was used for the partial layout conditioning examples in  Fig. 4, An overview is presented in  Algorithm 1. The idea is that at every timestep, the intermediate sample $\mathbf{x}_{i}(t)$ is first updated with the rate of change provided by our model ($v$). Then the sample is slightly adjusted to conform to a path which will yield the values of the partial conditioning layout $\mathbf{y}_{i}$ at non-null entries after inference. The adjustment per iteration is small, allowing the model to obey the conditioning and plausibly completing the rest of the layout.
To fully understand the algorithm, we provide additional details by first defining notation: The partial layout representation $\mathbf{y}_{i}=\{\mathbf{y}_{i}^{j}\}_{j\in J}$ is defined like the layout representation in  Sec. 3.2 extended by null values $\varnothing$, a placeholder value for entries of $\mathbf{y}_{i}$ tokens where no conditioning is provided. To give an example, consider a partial conditioning layout where the token $\mathbf{y}_{i}^{j}$ enforces that a bounding box with the label chair must be present in the final layout, but can have any coordinates. We set ${\mathbf{b}_{i}^{j}}=(\varnothing,\varnothing,\varnothing,\varnothing)$, ${\mathbf{c}_{i}^{j}}$ to be the PCA-reduced CLIP embedding of the word chair and $\alpha_{i}^{j}=1$, and write:
$\mathbf{y}_{i}^{j}=({\mathbf{b}_{i}^{j}}\mathbin{\|}{\mathbf{c}_{i}^{j}}\mathbin{\|}\alpha_{i}^{j})\,.$ (10)
The mask variable $M$ on line $5$ of our algorithm tracks which values of $\mathbf{y}_{i}$ are null-values, and masks these values out during the update on line $12$. To perform this masking, we define the arithmetic on $\varnothing$ as follows:
$\begin{array}[]{l}\varnothing+a=\varnothing\text{ for }a\in\mathbb{R},\\ \varnothing*a=\varnothing\text{ for }a\in\mathbb{R}-\{0\},\\ \varnothing*0=0.\end{array}$ (11)
Finally, we note that this algorithm is a generalization of the inference algorithm presented in  Sec. 3.3 of our main paper, which is the special case when no conditioning is provided ($\mathbf{y}_{i}\equiv\varnothing$)
0:  Input parameters description
0:  Output description
1:  conditionedInference( $P_{i}$ , $\mathbf{y}_{i}$ ):
2:  $T\leftarrow 1200$
3:  $\Delta t\leftarrow 1/T$
4:  $t\leftarrow 0$
5:  $M\leftarrow 0\text{ where }\mathbf{y}_{i}=\varnothing\text{ otherwise }1$ //Create a binary mask for the conditioning layout
6:  $\mathbf{x}_{i}(0)\sim\mathcal{N}(0,I)$ //Sample the starting noise
7:  while $t<1$ do
8:     $\frac{d\mathbf{x}_{i}(t)}{dt}\leftarrow v(\mathbf{x}_{i}(t),t,P_{i})$ //Calculate the rate of change of $\mathbf{x}_{i}(t)$ at timestep $t$
9:     $t\leftarrow t+\Delta t$ //Update timestep $t$
10:     $\mathbf{x}_{i}(t)\leftarrow\mathbf{x}_{i}(t-\Delta t)+\frac{d\mathbf{x}_{i}(t-\Delta t)}{dt}\cdot\Delta t$ //Calculate $\mathbf{x}_{i}(t)$ for the next timestep
11:     $\mathbf{y}_{i}(t)\leftarrow\mathbf{y}_{i}\cdot t+\mathbf{x}_{i}(0)\cdot(1-t)$ //Calculate conditioning update $\mathbf{y}_{i}(t)$
12:     $\mathbf{x}_{i}(t)\leftarrow\mathbf{y}_{i}(t)\odot M+\mathbf{x}_{i}(t)\odot(1-M)$ //Update $\mathbf{x}_{i}(t)$ with conditioning in masked area
13:  end while
14:  Return $\mathbf{x}_{i}(1)$
Algorithm 1 Partially Conditioned Layout Generation
Appendix C Training Data and Hyperparameters
Our model consists of $20$ AdaLN transformer blocks with $12$-headed attention. For a token $\mathbf{x}^{j}$, we sinusoidally encode $\mathbf{b}^{j}$ into $\mathbb{R}^{72}$, and $\alpha^{j}$ into $\mathbb{R}^{18}$. $\mathbf{c}^{j}$ consists of the $30$ top principal components of the object-label’s CLIP embedding, which accounts for $77.35\%$ of the explainable variance of our embeddings found in our training data. The timestep $t$ is sinusoidally encoded into $\mathbb{R}^{9}$, while the CLIP embedding of a global prompt $\ell$ is down-projected by a trainable linear layer into $\mathbb{R}^{17}$ before interfacing with the AdaLN block.
When reporting model parameters, we include all transformer block weights and attached linear layers, including the PCA projection matrices. Given that CLIP dominates the number of parameters, it is a necessary subcomponent for InstanceDiffusion, and needed to form any complete text-to-image pipeline, we factor it out.
We train our model for $2000$ epochs using stochastic gradient descent with learning rate $\lambda=0.0005$ and a batch size of $32$.
Appendix D Ablation of Image Generators
To control for the effect of the conditional generation component on our evaluation, we also run our text-to-image pipeline on competitors to InstanceDiffusion  [45], namely LMD+  [21], GliGEN  [20], and BoxDiff  [48], using the same image layouts we used in our main paper.
We show the generated image metrics for LMD+ in  Tab. 2. Like in our main paper, these metrics are inconclusive, and furthermore, the rankings of different layout generation methods differ depending on the downstream image generator.
LMD+
Model FID ($\downarrow$) KID ($10^{-2}$)($\downarrow$) CMMD ($\downarrow$) IS ($\uparrow$)
LayoutTransformer 0.730.730.730.730.73 $1.26$ $0.69$DeclareFontSeriesDefault[rm]bfb7.81
GPT4o $1.73$ 1.171.171.171.171.17 $0.69$ 7.667.667.667.667.66
Fixed Layout $3.81$ $1.93$ $0.77$ $7.06$
No Layout $1.82$ $1.87$ $0.83$ $7.33$
[…]
GliGEN
Model FID ($\downarrow$) KID ($10^{-2}$)($\downarrow$) CMMD ($\downarrow$) IS ($\uparrow$)
[…]
GPT4o $3.49$ $1.30$ $0.63$ 8.138.138.138.138.13
Fixed Layout $2.88$ $2.51$ $0.68$ $6.57$
No Layout $1.85$ $1.89$ $0.82$ $7.51$
[…]
BoxDiff
Model FID ($\downarrow$) KID ($10^{-2}$)($\downarrow$) CMMD ($\downarrow$) IS ($\uparrow$)
LayoutTransformer $1.37$ 1.621.621.621.621.62 $0.53$DeclareFontSeriesDefault[rm]bfb9.049.049.049.04
GPT4o 1.231.231.231.231.23 $1.77$ $0.63$ $7.40$
Fixed Layout $2.32$DeclareFontSeriesDefault[rm]bfb1.421.421.421.42 0.520.520.520.520.52 $6.92$
No Layout $1.82$ $1.87$ $0.83$ $7.33$
[…]
Table 2: Metrics evaluated on LMD+, GliGEN, and BoxDiff. For LMD+, our method is lowest in FID and KID. Different from the results on the InstanceDiffusion-generated images shown in Tab. 1, it is now also lowest in CMMD. For GliGEN, our method now performs second-best on FID and KID, but best on CMMD and IS. This is opposite to the results of InstanceDiffusion-generated images shown in Tab. 1 For BoxDiff, our method is lowest in FID, we are not the lowest in KID anymore. We now perform the best in CMMD and second best in IS, unlike in images generated from InstanceDiffusion.
We also run a human evaluation on the images from each of these three generators, shown in  Fig. 10. Although our results are not identical between image generation methods, there are several important commonalities across all four of these methods: our method is significantly more plausible than LayoutTransformer, and has significantly more variety than GPT4o and No Layout. We even outperform GPT4o in terms of plausibility for LMD+ and GliGEN. These results support that our method offers an effective balance between plausibility and variety, regardless of the choice of image generator. These also show that our human evaluation procedure can help control for the choice of image generator.
Refer to caption Refer to caption Refer to caption
Figure 10: Human evaluations of plausibility and variability of our pipeline using the LMD+, GliGEN, and BoxDiff generators. Across these evaluators, we see that our method has higher plausibility than LayoutTransformer, and higher variety than GPT4o and No Layout. In the case of LMD+ and GliGEN, we even have higher plausibility than GPT4o
Appendix E LayoutTransformer Temperature
Throughout our main paper, we adapted LayoutTransformer as faithfully as possible to the original work, meaning that we kept the temperature parameter equal to one. However, the question arises whether the generated layouts would be higher quality at lower temperatures, where the model’s output is more stable. Therefore, we examined the model’s performance when setting the temperature to the lowest possible setting of zero.
The results in  Tab. 3 show that even though layout object positions improve at the lowest temperature, although this comes at the cost of poorer object numeracy and a complete lack of variance among produced layouts for the same prompt. The changes in the robust traditional metrics were also inconclusive, with FID worsening and KID improving at the lowest temperature. Furthermore, LayoutTransformer with a temperature of zero still performed worse than our method on FID, KID, and our layout measurement metrics. In  Fig. 11, we show qualtatively that the final images produced at lower temperatures are still lacking in plausibility.
Even in this best-case scenario when we select the lowest temperature for optimal stability, we are still not measuring a decisive improvement across numerical metrics. Therefore, we saw no clear indication that setting the temperature parameter lower than the default of one would lead to missed improvements in model performance on the human evaluation, and kept the temperature at its original setting of one to remain as faithful as possible to the prior work.
Model FID ($\downarrow$) KID ($10^{-2}$)($\downarrow$) CMMD ($\downarrow$) IS ($\uparrow$) $O_{\text{Num}}(\downarrow)$ $l_{\text{Pos}}^{(1)}(10^{-11})(\uparrow)$ $l_{\text{Pos}}^{(2)}(10^{-11})(\uparrow)$ $\sigma^{2}_{\text{Pos}}(\uparrow)$
LayoutTransformer temp= 1 0.440.440.440.440.44 $0.94$DeclareFontSeriesDefault[rm]bfb1.341.341.341.34 $8.34$DeclareFontSeriesDefault[rm]bfb0.900.900.900.90 $3.09$ $1.21$DeclareFontSeriesDefault[rm]bfb231231231231
LayoutTransformer temp= 0 $0.48$ 0.920.920.920.920.92 $1.77$DeclareFontSeriesDefault[rm]bfb8.748.748.748.74 $4.11$ 3.733.733.733.733.73 1.531.531.531.531.53 00
[…]
Table 3: Comparison of metrics LayoutTransformer with a temperature of one (model default) and a temperature of zero. Even when the temperature is zero, we see that our method still performs better than the LayoutTransformer on the FID and KID metrics, which were shown to be robust in  Tab. 1. We also perform better than temp = $1$ in $O_{\text{Num}}$, $l_{\text{Pos}}^{(1)}$, $l_{\text{Pos}}^{(2)}$, and $\sigma^{2}_{\text{Pos}}$.
“Street” “Kitchen” “Bathroom”
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 11: Example layouts and images for LayoutTransformer when temp=00. Even at the most stable setting, the images appear implausible. Objects that are typically small details, such as bag, glass, or bottle repeated many times across the layout.
Appendix F GPT4o Temperature
Because we observed low image variance for GPT4o layouts, we also considered what would happen if we raised the temperature of GPT4o from the default $0.25$ as set in LLM-grounded Diffusion to achieve more variety.
We experimented with increasing the temperature from $0.2$ in increments of $0.1$. We found that at a temperature of one, GPT4o failed to produce a parsable layout $14\%$ of the time. However, these mistakes were easy to catch and query the model again. Temperatures higher than $1$ caused more frequent parsing failures, and began to produce long, tangential sentences rather than proper object labels. It was not possible to heuristically filter these responses, and so we settled on a temperature of one as a reasonable upper limit for operation temperature of GPT4o on this task.
We compare the performance of GPT4o with a temperature of one with our method, and GPT4o with the default temperature in  Tab. 4. Our model still outperforms GPT4o when the temperature is one in FID and KID. While raising the temperature improves the object numeracy score $O_{\text{Num}}$ and the positional variance score $\sigma^{2}_{\text{Pos}}$ improve in GPT4o when the temperature is raised, they are still worse than our method, and come at the cost of decreased performance in the positional likelihood scores $l_{\text{Pos}}^{(1)}$ and $l_{\text{Pos}}^{(2)}$. Therefore, raising the temperature does not offer a clear advantage on our numerical metrics.
We also visualized outputs of GPT4o with the raised temperature in  Fig. 12. Although there is some increase in the variation of scenes, the effect does not appear to be noticeably pronounced. Therefore, we choose to stick with a temperature of $0.25$ for our human evaluation, as this is the most faithful adaptation of our LLM-grounded Diffusion baseline, without neglecting a clear optimization.
Model FID ($\downarrow$) KID ($10^{-2}$)($\downarrow$) CMMD ($\downarrow$) IS ($\uparrow$) $O_{\text{Num}}(\downarrow)$ $l_{\text{Pos}}^{(1)}(10^{-11})(\uparrow)$ $l_{\text{Pos}}^{(2)}(10^{-11})(\uparrow)$ $\sigma^{2}_{\text{Pos}}(\uparrow)$
GPT4o temp=$0.25$ 0.940.940.940.940.94 $0.99$DeclareFontSeriesDefault[rm]bfb1.341.341.341.34 7.877.877.877.877.87 $3.71$DeclareFontSeriesDefault[rm]bfb4.374.374.374.37 1.491.491.491.491.49 $93$
GPT4o temp=1 $1.47$ $1.62$ $1.35$DeclareFontSeriesDefault[rm]bfb7.967.967.967.96 2.862.862.862.862.86 $4.02$ $1.35$ 142142142142142
[…]
Table 4: Comparison of metrics GPT4o with a temperature of $0.25$ (adapted model default) and one (highest stable temperature). At increased temperatures, GPT4o performs worse on the FID and KID metrics, which were shown to be robust in  Tab. 1. Although increasing the temperature of GPT4o improves $O_{\text{Num}}$ (the object frequencies are closer to the ground truth) and $\sigma^{2}_{\text{Pos}}$ (the layouts are more varied overall), performance on $l_{\text{Pos}}^{(1)}$ and $l_{\text{Pos}}^{(2)}$ drops (the positions of the objects are less plausible). Our method still performs better in $O_{\text{Num}}$, $l_{\text{Pos}}^{(1)}$, $l_{\text{Pos}}^{(2)}$, and $\sigma^{2}_{\text{Pos}}$ than GPT4o with a raised temperature.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 12: Example images and underlying layouts of the prompt street for GPT4o when the temperature is one, the highest stable temperature. Visually, there is slightly more variation than at a temperature of $0.25$, (see  Fig. 23), but this is not a pronounced effect: positions and quantities of objects, and implied camera angle, are still very repetitive.
Appendix G GPT4o Query Template
We adapted the prompt template from LLM-grounded Diffusion by replacing the old scene caption and layout examples with ground truth data from ADE20k and encourage chain of thought reasoning  [46]. Our LLM prompt is listed in  Fig. 13. Our in-context examples are listed in  LABEL:tab:incontext.
1 Task Description and Rules
2 You are a smart program for automatic image layout generation. I provide you with a global prompt which describes the entire image. The image layout has a height of 512 and a width of 512. The coordinate system assumes the origin (0,0) is in the top left corner. Bounding box coordinates are specified in the format (x,y,w,h), where x and y are the top left corner coordinate, and w and h are the full width and height of the box. Your task is to imagine which objects reasonably belong in an image with a global prompt, and arrange these objects in into a layout which could plausibly be for a real image.
3
4 Meta Command
5 Reason about the objects added to the layout For each object reason about its position in the layout relative to the other objects, and why it is likely. In general maintain a plausible configuration of the objects within the image layout such that the coordinates obey our coordinate convention. Do not number the objects, instead put them in a list in the exact format shown below. Remember to include the caption, background prompt and negative prompt in the layout.
6
7 [ In - context Examples ]
8
9 Question
10 Provide the layout for a "{prompt}"
Figure 13: Our full prompt to the LLM for layout generation. LLM starts completions from “Objects.”
Table 5: Our in-context examples. We use fixed in-context examples for layout generation.
1 Provide a layout for an "airport terminal”
2 Answer:
4 Caption: An airport terminal
5 Objects: [(’ceiling’,[1, 0, 510, 292]), (’floor’,[0, 360, 468, 151]), (’wall’,[0, 337, 152, 152]), (’wall’,[3, 193, 182, 100]),(’first floor’,[353, 262, 158, 96]),(’wall’,[353, 207, 158, 86]),(’first floor’,[0, 264, 136, 93]),(’plant’,[177, 398, 158, 78]),(’fountain’,[197, 376, 122, 79]),(’first floor’,[134, 310, 265, 30]),(’wall’,[398, 343, 68, 117]),(’wall’,[467, 330, 44, 172]),(’wall’,[190, 270, 157, 42]),(’column’,[65, 159, 30, 157]),(’column’,[0, 76, 19, 242]),(’column’,[442, 153, 30, 141]),(’wall’,[275, 336, 123, 33]),(’wall’,[134, 340, 129, 31]),(’fence’,[331, 394, 44, 83]),(’first floor’,[110, 282, 86, 39]),(’column’,[404, 171, 23, 139]),(’wall’,[239, 499, 272, 12]),(’column’,[113, 175, 22, 144]),(’seat’,[187, 478, 134, 17]),(’column’,[380, 193, 19, 117]),(’fence’,[134, 403, 40, 53]),(’column’,[145, 196, 18, 116]),(’tree’,[345, 324, 34, 45]),(’tree’,[43, 404, 30, 49]),(’tree’,[150, 329, 32, 45])]
6 Background prompt: an airport terminal
7 Negative prompt: empty
8 Reasoning: Airport terminals contain many walls and columns, and have a floor and ceiling. They also contain seats for passengers to wait in as well as decorative trees
9
10 Provide a layout for an "schoolhouse”
11 Answer:
12 Caption: schoolhouse
13 Objects: [(’sky’,[0, 0, 510, 431]),(’building’,[22, 23, 460, 465]),(’tree’,[1, 0, 173, 200]),(’grass’,[0, 449, 510, 61]),(’tree’,[422, 129, 83, 281]),(’tree’,[1, 202, 44, 254]),(’path’,[0, 478, 308, 27]),(’grass’,[0, 469, 201, 27]),(’plant’,[42, 412, 71, 76]),(’plant’,[399, 417, 56, 69]),(’person’,[229, 412, 33, 64]),(’car’,[3, 444, 40, 29]),(’tree’,[0, 431, 41, 23]),(’tree’,[472, 426, 37, 23])]
14 Background prompt: schoolhouse
15 Negative prompt: empty
16 Reasoning: A schoolhouse is typically a building. The layout could include a path, students, trees, plants, and a car in the schoolyard.
17
18 Provide a layout for an "ball pit”
19 Answer:
20 Caption: ball pit
21 Objects: [(’inflatable park’,[1, 0, 510, 510]),(’person’,[85, 42, 313, 398]),(’ball’,[451, 292, 48, 69]),(’ball’,[77, 253, 46, 61]),(’ball’,[416, 278, 40, 58]),(’ball’,[475, 265, 34, 68]),(’ball’,[371, 240, 39, 55]),(’ball’,[430, 246, 40, 47])]
22 Background prompt: ball pit
23 Negative prompt: empty
24 Reasoning: A ball pit is an inflatable park with balls and people. The layout could include a person playing in the ball pit and colorful balls scattered around the inflatable park.
25
26 Provide a layout for an "jail cell”
27 Answer:
28 Caption: jail cell
29 Objects: [(’bar’,[0, 0, 510, 512]),(’floor’,[24, 304, 390, 206]),(’wall’,[296, 16, 156, 482]),(’wall’,[72, 4, 232, 302]),(’bed’,[174, 256, 234, 196]),(’cell’,[462, 26, 48, 484]),(’wall’,[20, 4, 50, 458]),(’shelf’,[66, 48, 242, 20]),(’sink’,[152, 194, 40, 54])]
30 Background prompt: jail cell
31 Negative prompt: empty
32 Reasoning: A jail cell typically has bars, walls, a floor, and a bed. The layout could include a cell door, a shelf, and a sink.
33
34 Provide a layout for an "badlands”
35 Answer:
36 Caption: badlands
37 Objects: [(’earth’,[0, 199, 334, 267]),(’earth’,[58, 201, 453, 144]),(’hill’,[0, 106, 512, 118]),(’sky’,[0, 0, 512, 116]),(’earth’,[194, 334, 316, 177]),(’water’,[34, 218, 301, 128]),(’tree’,[0, 369, 236, 142]),(’rock’,[0, 381, 97, 83]),(’person’,[463, 273, 18, 71]),(’tripod’,[450, 289, 5, 38]),(’photo machine’,[449, 283, 8, 8])]
38 Background prompt: badlands
39 Negative prompt: empty
40 Reasoning: Badlands are characterized by eroded rock formations, so the layout could include earth, hills, rocks, and trees. The badlands may also have water, a person, a tripod, and a photo machine.
41
42 Provide a layout for an "art gallery”
43 Answer:
44 Caption: art gallery
45 Objects: [(’wall’,[224, 36, 287, 360]),(’floor’,[0, 323, 512, 188]),(’wall’,[0, 84, 226, 261]),(’ceiling’,[0, 0, 511, 112]),(’board’,[306, 153, 205, 140]),(’board’,[0, 170, 250, 102]),(’double door’,[251, 176, 55, 168]),(’grill’,[378, 260, 21, 91]),(’grill’,[338, 257, 20, 85]),(’vent’,[248, 22, 47, 19]),(’drawing’,[490, 196, 21, 40]),(’spotlight’,[453, 32, 17, 49]),(’drawing’,[8, 194, 18, 35]),(’spotlight’,[381, 54, 14, 43]),(’drawing’,[456, 241, 22, 25]),(’spotlight’,[279, 83, 15, 35]),(’spotlight’,[320, 71, 14, 38]),(’drawing’,[391, 187, 17, 30]),(’drawing’,[314, 201, 20, 26]),(’vent’,[6, 45, 36, 13]),(’spotlight’,[259, 88, 12, 34]),(’drawing’,[420, 196, 14, 28]),(’drawing’,[445, 204, 18, 22]),(’drawing’,[409, 239, 13, 28]),(’spotlight’,[234, 97, 11, 32]),(’drawing’,[43, 195, 18, 18]),(’drawing’,[351, 181, 12, 27]),(’drawing’,[135, 237, 17, 19]),(’drawing’,[40, 227, 14, 23]),(’drawing’,[205, 200, 11, 27])]
46 Background prompt: art gallery
47 Negative prompt: empty
48 Reasoning: An art gallery is indoors, so it has walls, a floor, and a ceiling. It can also have boards for displaying art, doors, grills, vents, and spotlights. The art gallery may have drawings on the walls and spotlights to illuminate the art.
49 Provide a layout for an "art gallery”
50 Answer:
51 Caption: window seat
52 Objects: [(’seat’,[2, 172, 507, 337]),(’floor’,[28, 322, 482, 187]),(’wall’,[102, 0, 266, 228]),(’wall’,[0, 0, 109, 510]),(’person’,[222, 20, 133, 390]),(’wall’,[363, 0, 146, 324]),(’windowpane’,[140, 0, 204, 69]),(’windowpane’,[0, 0, 102, 122]),(’windowpane’,[388, 0, 122, 75]),(’hat’,[375, 157, 80, 69])]
53 Background prompt: window seat
54 Negative prompt: empty
59 Reasoning: A window seat typically has a seat, walls, and a floor. The layout could include a person sitting on the seat, looking out the window, and wearing a hat.
Appendix H Comparison to DDIM
We initially considered a DDIM  [43] based approach rather than rectified flow. However, early experiments showed less promise in this direction. DDIM models struggled with generating the correct CLIP embeddings, leading to meaningless images that did not match the prompt, whereas rectified flow-based approaches were more successful without needing to search the hyperparameter space.
We provide an example here, from a model with an identical architecture to our presented model (including all hyperparameters specified in  Appendix C, except it is trained with a DDIM training objective and performs DDIM inference (with a log-linear noise schedule from $\sigma=0.02$ to $\sigma=1$). This is not an exhaustive search by any means, but is intended as a point-of-reference for other researchers.
We show our statistics in  Tab. 6, and some visual examples from the model in  Fig. 14. This model performs worse than ours for numerical metrics, which were robust against adversarial examples in  Sec. 5.5.
We are not entirely sure why we observe this performance between the two techniques. We speculate that the straighter transit paths of samples rectified flow  [24] increases the model’s ability to effectively learn high dimensional data like the PCA-reduced CLIP embeddings. However, this will require more investigation to confirm.
Model FID ($\downarrow$) KID ($10^{-2}$)($\downarrow$) CMMD ($\downarrow$) IS ($\uparrow$) $O_{\text{Num}}(\downarrow)$ $l_{\text{Pos}}^{(1)}(10^{-11})(\uparrow)$ $l_{\text{Pos}}^{(2)}(10^{-11})(\uparrow)$ $\sigma^{2}_{\text{Pos}}(\uparrow)$
Ours (DDIM) $0.95$ $8.60$ $1.77$DeclareFontSeriesDefault[rm]bfb7.897.897.897.89 $4.33$ $0.25$ $0.01$DeclareFontSeriesDefault[rm]bfb239239239239
[…]
Table 6: Generated image metrics, and our generated layout numerical metrics applied on our model architecture with DDIM or rectified flow. Our model performs better on image metrics that were robust to adversarial examples in  Tab. 1, i.e., FID and KID. DDIM performs poorly on $O_{\text{Num}}$, corroborating our observations of improper object labels in  Fig. 14, which also is responsible for the poor scores in $l_{\text{Pos}}^{(1)}$ and $l_{\text{Pos}}^{(2)}$. Although $\sigma^{2}_{\text{Pos}}$ is higher for DDIM, this is at the cost of the layouts being largely nonsense (see  Fig. 14)
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 14: Our Model with DDIM instead of Rectified Flow - Street. The bounding box labels match poorly to the desired scene, and the resulting images appear to be implausible.
Appendix I Additional Images and Layouts
Here we present additional examples of our model’s generated layouts, and conditionally generated images, for the prompts bedroom (Fig. 15), mountain (Fig. 16), and kitchen (Fig. 17).
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 15: Ours - Bedroom.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 16: Ours - Mountain.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 17: Ours - Kitchen.
Appendix J Print-ready Main Results Diagram
For readers who prefer the document on paper, we include our visual results diagram from  Fig. 3 at a size where the annotations are large enough to be printed clearly. The annotations can also be zoomed into on our main document.
Refer to caption Refer to caption Refer to caption
Figure 18: No Layout - Living Room.
Refer to caption Refer to caption Refer to caption
Figure 19: No Layout - Roof Top
Refer to caption Refer to caption Refer to caption
Figure 20: No Layout - Street
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 21: GPT4o - Living Room.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 22: GPT4o - Roof Top.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 23: GPT4o - Street.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 24: Layout Transformer - Living Room.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 25: Layout Transformer - Roof Top.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 26: Layout Transformer - Street.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 27: Ours - Living Room.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 28: Ours - Roof Top.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 29: Ours - Street.